# The nuclear-spin dark state in silicon

### v1

@@ Abstract
We report the formation of a nuclear-spin dark state in a gate-defined silicon double quantum dot. The dark state has been predicted to emerge during dynamic nuclear polarization when the electrons in the quantum dot drive the nuclei in the semiconductor into a decoupled state. We show that, as expected, the electron-nuclear coupling rapidly diminishes in the dark state, and that this state depends on the synchronized precession of the nuclear spins. Moreover, the dark state significantly increases the lifetimes of electronic spin states. This nuclear-spin dark state could be used as a quantum memory or as a resource for quantum sensing, and further studies of the dark state may point the way toward increased polarization of nuclear spin ensembles.
@@ I Introduction
Nuclear spins are abundant in condensed matter systems, and the ability to manipulate large ensembles of nuclear spins has led to many technological advances. Nuclear magnetic resonance, for example, has profoundly shaped science and medicine. In the quantum realm, nuclear spin ensembles often create decoherence for semiconductor qubits based on electron spins, and the ability to precisely control these nuclei would be beneficial for applications like quantum computing. Indeed, many ideas from nuclear magnetic resonance have been employed to mitigate this decoherence. One such approach, dynamic nuclear polarization (DNP), aims to suppress fluctuations in nuclear spin ensembles by fully polarizing the nuclei. Platforms including bulk solids [1, 2], gate-defined quantum dots [3, 4, 5], and self-assembled quantum dots [6, 7] all offer routes to DNP, but achieving full polarization remains a challenge. Significant theoretical research has focused on understanding why this is the case. Multiple lines of research have predicted that a collective nuclear-spin dark state, which does not interact with electron spins, should emerge during DNP  [8, 9, 10, 11, 12, 13, 14, 15]. Understanding the nuclear-spin dark state is crucial to improving DNP, and the dark state is also predicted to be an excellent quantum memory and resource for quantum sensing [9, 16, 17].
@@
Here, we present evidence for the formation of the nuclear-spin dark state in a Si gate-defined quantum dot. We directly probe the electron-nuclear coupling through Landau-Zener sweeps, and we find that it sharply reduces during DNP, in agreement with theoretical predictions. We also observe that the dark state persists for about 1 ms, in agreement with nuclear spin inhomogeneous dephasing times in Si. We also verify that the dark state significantly increases the lifetime of the electronic spin states involved. Together with these experiments, we construct numerical simulations that model the quantum-mechanical behavior of the electrons, as well as the semi-classical behavior of each of the roughly $10^{4}$ nuclear spins interacting with the electrons in our device. Our work verifies key theoretical predictions about the nuclear-spin dark state. Moreover, our work demonstrates that exotic nuclear spin states can occur in the very same devices that are used for quantum computing [20, 21, 22], potentially opening the door to using such states as quantum memories or as resources for quantum sensing or suppressing decoherence.
@@ II Setup
We operate the device near the $(4,0)$$(3,1)$ charge transition, with 4(0) or 3(1) electrons in dot 1(2). In this configuration, two electrons in dot 1 always fill the lowest valley level as a singlet. The system dynamics can thus be described within a two-electron picture for the remaining “valence” electrons, which can form an effective singlet-triplet qubit [24, 25]. Figure 1(b) shows the relevant energies of spin states in the $(3,1)$ charge configuration as a function of the detuning $\epsilon$. At a specific detuning $\epsilon_{ST}$, the exchange coupling $J$ equals the electronic Zeeman energy $E_{Z}=\bar{g}\mu_{B}B^{z}_{tot}$, where $\bar{g}\approx 2$ is the electron $g$-factor in Si, and $B^{z}_{tot}$ is the total magnetic field experienced by the electrons, which includes both the externally applied field $B^{z}$ and the longitudinal hyperfine field. At $\epsilon_{ST}$, the spin singlet state, $|S\rangle=(\mid\uparrow\downarrow\rangle-\mid\downarrow\uparrow\rangle)/\sqrt{2}$, comes into resonance with the polarized triplet $|T_{-}\rangle=\mid\downarrow\downarrow\rangle$. The hyperfine interaction between the electron and nuclear spins couples these states. In particular, a difference in the transverse hyperfine field between the two dots couples $S$ and $T_{-}$, and the absolute value of this coupling is denoted by $\Delta_{ST}$. Deep in the (3,1) charge configuration, a difference in the longitudinal hyperfine field couples $S$ and $T_{0}$ [26].
@@ III Results
Figure 1(c) shows the results of an experiment with the external magnetic field $B^{z}=1$ mT. We repeatedly initialize the electrons as a singlet and then pulse $\epsilon$ to different values near $\epsilon_{ST}$ for a fixed time of $4~{}\mu$s before the measurement. In the left panel, the thin line indicates the points at which the $S\text{-}T_{-}$ transition occurs [27]. We observe that $\epsilon_{ST}$ increases over time, corresponding to a reduction in $B^{z}_{tot}$. This change in the total field is a hallmark of DNP in double quantum dots and has been observed previously in GaAs double quantum dots [4, 5]. As described in the Supplemental Material, we also observe that the longitudinal magnetic Zeeman gradient $\Delta E_{Z}$ changes during a DNP process, consistent with results from GaAs devices. We do not observe significant DNP effects at larger magnetic fields. This likely occurs because spin-orbit effects [28, 27], which can quench DNP, become significant at higher fields.
@@
Curiously, the $S\text{-}T_{-}$ transition probability appears to decrease with time. This is revealed by the decreasing signal amplitude at $\epsilon_{ST}$ over time [Fig. 1(c), right panel]. To explore this feature with better control, we perform DNP via Landau-Zener (LZ) sweeps through $\epsilon_{ST}$ [4, 5], [Fig. 1(b)]. We measure how the LZ transition probability depends on the number of sweeps. We repeat a “pump” cycle, consisting of a singlet initialization, an LZ sweep through $\epsilon_{ST}$, and a single-shot singlet-triplet measurement, 64 times [Fig. 2(a)]. During the LZ sweep, we sweep $\epsilon$ from 6.2 to 7.7 mV through $\epsilon_{ST}$ in $5$ $\mu$s, with a time interval $\tau_{w1}$ of $27~{}\mu$s to the next sweep. After this set of 64 pump cycles, we halt the pumping for 5.12 ms to allow the nuclear spin system to dephase and to implement auxiliary pulses to calibrate our measurement (see Supplemental Material). Figure 2(b) shows the triplet probability $\langle\it{P}_{T}\rangle$ after each LZ sweep as a function of the sweep number, averaged over more than $10^{4}$ pump-pause cycles.
@@
The experimental data reveal that $\langle P_{T}\rangle$ rapidly decreases with the sweep number and saturates at a small value [Fig.2(b)]. The remainder of this paper will explore this phenomenon. One possible explanation for this effect is that our singlet-triplet readout contrast diminishes with increased nuclear polarization. Such an effect has been observed in past work in GaAs quantum dots [29, 30]. This effect occurs because the $T_{0}$ state relaxation rate increases significantly when the longitudinal magnetic gradient $\Delta E_{Z}$ becomes large, usually on the order of hundreds of MHz. However, this effect is unlikely to explain our data, because the typical $\Delta E_{Z}$ values are on the order of MHz in our device. In addition, we measure the $T_{-}$ state, which does not experience increased relaxation at large $\Delta E_{Z}$. Another possible explanation for our data is that the process of DNP is so efficient that $\epsilon_{ST}$ moves outside of our LZ sweep window. As discussed further below, and as shown in Fig. 4, this is also unlikely.
@@
A third explanation is the formation of the nuclear-spin dark state. The dark state should emerge during DNP as the nuclei are driven into a state that decouples from the electrons [10, 14]. In the case of a double quantum dot, the dark state is characterized by a vanishing net transverse hyperfine gradient ($2\Delta_{ST}$ in magnitude). Figure 2(b) shows the root-mean-square values of $\Delta_{ST}$ extracted from experimental data, using the analytical expression for the LZ transition probability [31, 32] as $\langle P_{T}(t)\rangle=\langle 1-\exp[-\frac{(2\pi)^{2}\Delta_{ST}^{2}(t)}{\beta}]\rangle\approx\frac{(2\pi)^{2}\langle\Delta_{ST}^{2}(t)\rangle}{\beta}$, where $\beta$ is the sweep rate [Table 1]. $\langle\Delta_{ST}^{2}\rangle^{1/2}$ rapidly decreases with the LZ sweep number from 110 to 30 KHz.
@@
The dark state can be understood in the following semiclassical picture. In thermal equilibrium, the average polarization along any direction of an ensemble of $N$ nuclear spins is approximately $\sqrt{N}$. Thus, the average imbalance in the transverse polarization between two dots with separate nuclear spin ensembles is also of order $\sqrt{N}$. During an LZ sweep, the quantum spin state of the electrons creates an effective magnetic field (the Knight field) that causes the difference of the transverse nuclear polarizations to rotate out of the $x$-$y$ plane and toward the $z$ direction. Thus, after about $\sqrt{N}$ LZ sweeps, the difference in the transverse polarization should vanish, leading to a scenario where $\Delta_{ST}=0$. In this “dark state”, the transverse polarizations of the two dots are the same, and they undergo a synchronized precession around the external field.
@@
We perform numerical simulations of the full electron-nuclear system using the semiclassical techniques described in Refs. [10, 12] and the parameters in Table 1. This approach combines a quantum description of the two electron spins with a semiclassical model of the coherent dynamics for each of the nuclear spins interacting with the electrons in the quantum dots. At each time during the LZ sweeps, we compute the quantum state of the electrons, the effective Knight field exerted by the electrons on the nuclear spins, and how the nuclei evolve in this Knight field. We account for nuclear spin dephasing and relaxation. A nuclear spin relaxation time of $T_{1,n}=5$ s is chosen to reproduce the saturation $\langle P_{T}\rangle$ value in Fig. 2(b), and is broadly consistent with other measurements we have made. We also account for electrical noise experienced by the electrons. We consider 8000 and 4000 nuclei for dot 1 and dot 2, respectively, estimated according to the physical properties of the Si quantum dots. We use an hyperfine interaction strength of $|\bar{A}|=0.052$ GHz, corresponding to a two-electron spin dephasing time of approximately 0.8 $\mu$s [33]. The negative prefactor of $\bar{A}$ is determined by the signs of the electron and nuclear $g$-factors in Si. See the Supplemental Material for further details on our simulations.
@@
Figure 2(c) shows a simulated evolution of the nuclear spins during repeated LZ sweeps starting from a random state, confirming the semiclassical picture for the dark state formation. Figure 2(b) shows simulations of our experiment, which agree quantitatively with our measurements. The residual $\Delta_{ST}$ is associated with the finite time interval $\tau_{w1}$ between LZ sweeps (see Supplemental Material). One surprising feature of these data, which is captured by the simulations, is that $\Delta_{ST}$ reduces over about 10 sweeps, while the expected number of sweeps to reach the dark state is about $\sqrt{10^{4}}=10^{2}$, as shown in [Fig. 2(c)]. The explanation for this effect is that over the roughly $10^{4}$ pump-pause cycles (each of which contains 64 LZ sweeps), a significant longitudinal nuclear polarization builds up, effectively reducing the transverse polarization that reappears during the pause periods. The reduction of the transition probability in a number of sweeps reflecting the difference in transverse polarizations between the dots is a key feature of the nuclear-spin dark state.
We next investigate what happens during the pause periods. The semiclassical picture of the dark state provided above suggests that the synchronized precession of the net polarizations in each dot underlies vanishing electron-nuclear coupling. In this case, the primary mechanism for the nuclear spin system to exit the dark state involves inhomogeneous nuclear spin dephasing. The key idea is that after the characteristic dephasing time, a new, random transverse polarization of magnitude $\sqrt{N}$ appears in each dot. One might intuitively expect that nuclear spin relaxation should spoil the dark state. While the nuclear spin $T_{1}$ time does limit the lifetime of the dark state, dephasing occurs more rapidly.
@@
To model these data, we consider the evolution of $\Delta_{ST}$ under the influence of nuclear spin dephasing. Assuming that at $t_{n}=0$, $\Delta_{ST}=0$, we analytically find that $\langle\Delta_{ST}^{2}\rangle\propto 1-\exp[-(t_{n}/T_{c})^{\gamma}]$. The characteristic time is closely related to the nuclear spin dephasing time, given by $T_{c}=T^{*}_{2,n}/\alpha$, with a factor $\alpha$ difference. Assuming that each nuclear spin experiences independent, Gaussian magnetic noise with a white power spectrum, $\langle\Delta_{ST}^{2}\rangle$ follows an exponential function with $\gamma=1$ and $\alpha=2$. For quasi-static noise, $\langle\Delta_{ST}^{2}\rangle$ follows a Gaussian function with $\gamma=2$ and $\alpha=\sqrt{2}$. By incorporating the form of $\langle\Delta_{ST}^{2}\rangle$ into the expression for the LZ transition probability $\langle P_{T}(t)\rangle\approx\frac{(2\pi)^{2}\langle\Delta_{ST}^{2}(t)\rangle}{\beta}$, we fit the experimental data. In the case of white noise [Fig. 3(b)], the best-fit value of $T^{*}_{2,n}$ is 4.1 ms, while for quasistatic noise [Fig. 3(c)], the best-fit value is 2.4 ms. In the figures, we also show numerical simulations with the fitted $T^{*}_{2,n}$ values. White noise appears to better describe the experimental data [Fig. 3(b)].
@@
Figure 4: Increasing the electronic lifetime. (a) Pulse sequence for measurement with and without pumping. In the case without pumping, the LZ sweep in the pump cycles is replaced by a wait pulse at $\epsilon=0$. For the probe cycles, we initialize the singlet state and pulse to different values of $\epsilon$ for a variable evolution time $t_{e}$. (b) Measured $\langle P_{T}\rangle$ for $\epsilon$ values near $\epsilon_{ST}$ in the absence of pumping. (c) Measured $\langle P_{T}\rangle$ for $\epsilon$ values near $\epsilon_{ST}$ with pumping. (d) $\langle P_{T}\rangle$ at $\epsilon_{ST}$ as a function of $t_{e}$. The experimental data (filled triangles and circles) are extracted from (b) and (c), respectively, corresponding to the $\langle P_{T}\rangle$ maximum for each $t_{e}$ value. Numerical simulations for both cases are presented.
@@
One major potential benefit of the nuclear-spin dark state is the possibility of increased electron-spin lifetimes, due to the reduced hyperfine coupling. To test this, we measure the relaxation time of the singlet state near $\epsilon_{ST}$ with and without DNP. In the former case, the pulse sequence starts with 64 pump cycles to form the dark state, followed by a “probe” cycle, as shown in Fig. 4(a). In the case without pumping, we use an identical pulse sequence but replace each LZ sweep in the pump cycle by a wait at $\epsilon=0$ mV, away from $\epsilon_{ST}$. In each probe cycle, with the electronic state initialized as a singlet, we pulse $\epsilon$ abruptly to a value near $\epsilon_{ST}$, wait at that $\epsilon$ for a variable time $t_{e}$, allowing the state to relax, and then abruptly pulse $\epsilon$ back to 0 mV for the measurement. Again, each probability value is obtained from over $10^{3}$ repeated measurements.
@@
Figures 4(b) and (c) show the measured $\langle P_{T}\rangle$ after the wait near $\epsilon_{ST}$ for variable time $t_{e}$ without and with pumping, respectively. Without pumping, $\langle P_{T}\rangle$ at $\epsilon_{ST}$ increases over a time-scale of about a microsecond and saturates between 0.1 and 0.2, as indicated by the yellow region in Fig. 4(b) and filled triangles in Fig. 4(d). The relatively low saturation value reflects both the relatively large voltage resolution used in this measurement, as well as the presence of detuning noise in our system.
@@
With pumping, however, $\langle P_{T}\rangle$ remains nearly zero up to at least 10 $\mu$s [Fig. 4(d)], indicating strongly suppressed electronic-spin relaxation. In principle, $\langle P_{T}\rangle$ should remain small for up to the lifetime of the dark state. Future experiments will focus on verifying this lifetime enhancement. As above, our numerical simulations agree well with our data [Fig. 4(d)]. We also note that even in the dark state, Fig. 4(c) shows that $\epsilon_{ST}$ still falls within our LZ sweep window, as discussed above.

### v2

# The formation of a nuclear-spin dark state in silicon

@@ Abstract
Silicon-based qubits are often made by trapping individual electrons in quantum dots defined by electric gates. Quantum information can then be stored using the spin states of the electrons. However, the nuclei of the surrounding atoms also have spin degrees of freedom that couple to the electron spin qubits and cause decoherence. The emergence of a nuclear-spin dark state has been predicted to reduce this coupling during dynamic nuclear polarization, when the electrons in the quantum dot drive the nuclei in the semiconductor into a decoupled state. Here, we report the formation of a nuclear-spin dark state in a gate-defined silicon double quantum dot. We show that, as expected, the transverse electron-nuclear coupling rapidly diminishes in the dark state, and that this state depends on the synchronized precession of the nuclear spins. Moreover, the dark state significantly reduces the relaxation rate between the singlet and triplet electronic spin states. This nuclear-spin dark state has potential applications as a quantum memory or in quantum sensing, and might enable increased polarization of nuclear spin ensembles.
@@ I Introduction
Nuclear spins are abundant in condensed matter systems, and the ability to manipulate large ensembles of nuclear spins has led to many technological advances. Nuclear magnetic resonance, for example, has profoundly shaped science and medicine. In the quantum realm, nuclear spin ensembles often create decoherence for semiconductor qubits based on electron spins, and the ability to precisely control these nuclei would be beneficial for applications like quantum computing. Indeed, many ideas from nuclear magnetic resonance have been employed to mitigate this decoherence. One such approach, dynamic nuclear polarization (DNP), aims to suppress fluctuations in nuclear spin ensembles by fully polarizing the nuclei. Platforms including bulk solids [1, 2], gate-defined quantum dots [3, 4, 5], and self-assembled quantum dots [6, 7] all offer routes to DNP, but achieving full polarization remains a challenge. Significant theoretical research has focused on understanding why this is the case. Multiple lines of research have predicted that a collective nuclear-spin dark state, which has suppressed hyperfine coupling to the electrons, should emerge during DNP  [8, 9, 10, 11, 12, 13, 14, 15]. Understanding the nuclear-spin dark state is crucial to improving DNP, and the dark state is also predicted to be an excellent quantum memory and resource for quantum sensing [9, 16, 17].
@@
Here, we present evidence for the formation of the nuclear-spin dark state in a Si gate-defined quantum dot. We directly probe the hyperfine coupling between two electronic spin states (the singlet and polarized triplet), through Landau-Zener sweeps, and we find that it sharply reduces during DNP, in agreement with theoretical predictions. We also observe that the dark state persists for about 1 ms, in agreement with nuclear spin inhomogeneous dephasing times in Si. We also verify that the dark state significantly reduces singlet-triplet relaxation. Together with these experiments, we construct numerical simulations that model the quantum-mechanical behavior of the electrons, as well as the semi-classical behavior of each of the roughly $10^{4}$ nuclear spins interacting with the electrons in our device.
Our work verifies key properties of the nuclear-spin dark state, which has been predicted to appear in multiple systems, including gate-defined and self-assembled quantum dots, as well as solid-state materials. The ability to probe the dark state will improve our understanding of this collective state. In the future, this understanding could lead to increased levels of nuclear-spin polarization, which could in turn lead to increased electron-spin coherence times. Moreover, our work demonstrates that exotic nuclear spin states can occur in the very same devices that are used for quantum computing [20, 21, 22], potentially opening the door to using such states as quantum memories or as resources for quantum sensing.
@@ II Setup
We operate the device near the $(4,0)$$(3,1)$ charge transition, with 4(0) or 3(1) electrons in dot 1(2). In this configuration, two electrons in dot 1 always fill the lowest valley level as a singlet. The system dynamics can thus be described within a two-electron picture for the remaining “valence” electrons, which can form an effective singlet-triplet qubit [24, 25, 26]. We operate the device with four electrons, instead of the usual two for a singlet-triplet qubit, because using four electrons facilitates spin-blockade readout, even in the presence of low valley splittings [25, 27, 28, 23].
Figure 1(b) shows the relevant energies of spin states in the $(3,1)$ charge configuration as a function of the detuning $\epsilon$. At a specific detuning $\epsilon_{ST}$, the exchange coupling $J$ equals the electronic Zeeman energy $E_{Z}=\bar{g}\mu_{B}B^{z}_{tot}$, where $\bar{g}\approx 2$ is the electron $g$-factor in Si, and $B^{z}_{tot}$ is the total magnetic field experienced by the electrons, which includes both the externally applied field $B^{z}$ and the longitudinal hyperfine field. At $\epsilon_{ST}$, the spin singlet state, $|S\rangle=(\mid\uparrow\downarrow\rangle-\mid\downarrow\uparrow\rangle)/\sqrt{2}$, comes into resonance with the polarized triplet $|T_{-}\rangle=\mid\downarrow\downarrow\rangle$. The hyperfine interaction between the electron and nuclear spins couples these states. In particular, a difference in the transverse hyperfine field between the two dots couples $S$ and $T_{-}$ states, and the absolute value of this coupling is denoted by $\Delta_{ST}$ [29]. A transverse hyperfine field couples $S$ and $T_{-}$, because these states differ in the $z$-component of their spin angular momentum. In general, a transverse magnetic field is thus required to transition from one state to the other. However, a uniform transverse field would flip the spins in both dots, leading to no coupling between these states. A difference in the transverse field is required to preferentially flip one spin, leading to a transition from $S$ to $T_{-}$. Deep in the (3,1) charge configuration, a difference in the longitudinal hyperfine field couples $S$ and $T_{0}$ [29].
@@ III Results
Figure 1(c) shows the results of an experiment with the external magnetic field $B^{z}=1$ mT. We repeatedly initialize the electrons as a singlet and then pulse $\epsilon$ to different values near $\epsilon_{ST}$ for a fixed time of $4~{}\mu$s before the measurement. In the left panel, the thin line indicates the points at which the $S\text{-}T_{-}$ transition occurs [30]. We observe that $\epsilon_{ST}$ increases over time, corresponding to a reduction in $B^{z}_{tot}$. We have verified that this shift in $\epsilon_{ST}$ does not result from drifts in our superconducting magnet. This change in the total field is a hallmark of DNP in double quantum dots and has been observed previously in GaAs double quantum dots [4, 5]. As described in Methods, we also observe that the longitudinal magnetic Zeeman gradient $\Delta E_{Z}$ changes during a DNP process, consistent with results from GaAs devices. We do not observe significant DNP effects at larger magnetic fields. This likely occurs because spin-orbit effects [31, 30], which can quench DNP, become significant at higher fields.
@@
Curiously, the $S\text{-}T_{-}$ transition probability appears to decrease with time. This is revealed by the decreasing signal amplitude at $\epsilon_{ST}$ over time [Fig. 1(c), right panel]. To explore this feature with better control, we perform DNP via Landau-Zener (LZ) sweeps through $\epsilon_{ST}$ [4, 5], [Fig. 1(b)]. We measure how the LZ transition probability depends on the number of sweeps. We repeat a “pump” cycle, consisting of a singlet initialization, an LZ sweep through $\epsilon_{ST}$, and a single-shot singlet-triplet measurement, 64 times [Fig. 2(a)]. During the LZ sweep, we sweep $\epsilon$ from 6.2 to 7.7 mV through $\epsilon_{ST}$ in $5$ $\mu$s, with a time interval $\tau_{w1}$ of $27~{}\mu$s to the next sweep. After this set of 64 pump cycles, we halt the pumping for $\tau_{w}=5.12$ ms to allow the nuclear spin system to dephase and to implement auxiliary pulses to calibrate our measurement (see Methods). Figure 2(b) shows the triplet probability $\langle\it{P}_{T}\rangle$ after each LZ sweep as a function of the sweep number, averaged over more than $10^{4}$ pump-pause cycles.
@@
The experimental data reveal that $\langle P_{T}\rangle$ rapidly decreases with the sweep number and saturates at a small value [Fig.2(b)]. The remainder of this paper will explore this phenomenon. One possible explanation for this effect is that our singlet-triplet readout contrast diminishes with increased nuclear polarization. Such an effect has been observed in past work in GaAs quantum dots [32, 33]. This effect occurs because the $T_{0}$ state relaxation rate increases significantly when the longitudinal magnetic gradient $\Delta E_{Z}$ becomes large, usually on the order of hundreds of MHz. However, this effect is unlikely to explain our data, because the typical $\Delta E_{Z}$ values are on the order of MHz in our device (see Extended Data Fig. 1). In addition, we measure the $T_{-}$ state, which does not experience increased relaxation at large $\Delta E_{Z}$. Another possible explanation for our data is that the process of DNP is so efficient that $\epsilon_{ST}$ moves outside of our LZ sweep window. As discussed further below, and as shown in Fig. 4, this is also unlikely.
@@
A third explanation is the formation of the nuclear-spin dark state. The dark state should emerge during DNP as the nuclei are driven into a state with significantly reduced $S\text{-}T_{-}$ coupling [10, 14]. In the case of a double quantum dot, the dark state is characterized by a vanishing net transverse hyperfine gradient ($2\Delta_{ST}$ in magnitude). Figure 2(b) shows the root-mean-square values of $\Delta_{ST}$ extracted from experimental data, using the analytical expression for the LZ transition probability [34, 35] as $\langle P_{T}(t)\rangle=\langle 1-\exp[-\frac{(2\pi)^{2}\Delta_{ST}^{2}(t)}{\beta}]\rangle\approx\frac{(2\pi)^{2}\langle\Delta_{ST}^{2}(t)\rangle}{\beta}$, where $\beta$ is the sweep rate [Table 1]. $\langle\Delta_{ST}^{2}\rangle^{1/2}$ rapidly decreases with the LZ sweep number from 110 to 30 KHz.
@@
The dark state can be understood in the following semiclassical picture. In thermal equilibrium, the average polarization along any direction of an ensemble of $N$ nuclear spins is approximately $\sqrt{N}$. Thus, the average imbalance in the transverse polarization between two dots with separate nuclear spin ensembles is also of order $\sqrt{N}$. During an LZ sweep, the quantum spin state of the electrons creates an effective magnetic field (the Knight field) that causes the difference of the transverse nuclear polarizations to rotate out of the $x$-$y$ plane and toward the $z$ direction. The reason for this behavior is the following. In brief, during an LZ sweep, the electronic spin state changes to the extent that a transverse difference field exists. If the electronic spin-state changes, the Knight field will acquire a transverse component. This transverse component will tend to rotate the nuclear polarizations from the $x$-$y$ plane to the $z$ axis. The transverse component of the Knight field has different signs for the two dots (see the Methods), so the difference in the transverse polarizations between the dots will be effectively rotated out of the plane. Thus, after about $\sqrt{N}$ LZ sweeps, the difference in the transverse polarization should vanish, leading to a scenario where $\Delta_{ST}=0$. In this “dark state”, the transverse polarizations of the two dots are the same, and they undergo a synchronized precession around the external field.
@@
We perform numerical simulations of the full electron-nuclear system using the semiclassical techniques described in Refs. [10, 12] and the parameters in Table 1. This approach combines a quantum description of the two electron spins with a semiclassical model of the coherent dynamics for each of the nuclear spins interacting with the electrons in the quantum dots. At each time during the LZ sweeps, we compute the quantum state of the electrons, the effective Knight field exerted by the electrons on the nuclear spins, and how the nuclei evolve in this Knight field. We account for nuclear spin dephasing and relaxation. A nuclear spin relaxation time of $T_{1,n}=5$ s is chosen to reproduce the saturation $\langle P_{T}\rangle$ value in Fig. 2(b), and is broadly consistent with other measurements we have made (see Extended Data Fig. 1). We also account for electrical noise experienced by the electrons. We consider 8000 and 4000 nuclei for dot 1 and dot 2, respectively, estimated according to the physical properties of the Si quantum dots. We use an hyperfine interaction strength of $\bar{A}=-0.052$ GHz, corresponding to a two-electron spin dephasing time of approximately 0.8 $\mu$s [36]. The negative sign of $\bar{A}$ is determined by the signs of the electron and nuclear $g$-factors in Si. See Methods for further details on our simulations.
@@
Figure 2(c) shows a simulated evolution of the nuclear spins during repeated LZ sweeps starting from a random state, confirming the semiclassical picture for the dark state formation. In particular, the difference in transverse polarization decreases in about $\sqrt{10^{4}}=10^{2}$ sweeps, consistent with our expectation that $\sqrt{N}$ sweeps are required to reach the dark state. Figure 2(b) shows simulations of our experiment, which agree quantitatively with our measurements. The residual $\Delta_{ST}$ is associated with the finite time interval $\tau_{w1}$ between LZ sweeps (see Methods). One surprising feature of these data, which is captured by the simulations, is that $\Delta_{ST}$ reduces over about 10 sweeps, while the expected number of sweeps to reach the dark state is about $10^{2}$, as shown in Fig. 2(c). The explanation for this effect is that over the roughly $10^{4}$ pump-pause cycles (each of which contains 64 LZ sweeps), a significant longitudinal nuclear polarization builds up. As the longitudinal polarization increases, the torque from the Knight field during LZ sweeps can more effectively rotate the transverse polarization to the $z$-direction (see Methods), and fewer sweeps are required to reach the dark state.
We next investigate what happens during the pause periods. The semiclassical picture of the dark state provided above suggests that the synchronized precession of the net polarizations in each dot underlies the vanishing electron-nuclear coupling. In this case, the primary mechanism for the nuclear spin system to exit the dark state involves inhomogeneous nuclear spin dephasing. One might intuitively expect that nuclear spin relaxation should spoil the dark state. While the nuclear spin $T_{1}$ time does limit the lifetime of the dark state, dephasing occurs more rapidly. Within this picture, and starting from a random initial state, each dot has a net transverse nuclear polarization of order $\sqrt{N}$. Once the system enters the dark state after the LZ sweeps, the net transverse polarizations of the two dots have the same magnitude and phase. After the pause period, the net transverse polarizations of the dots have changed, because all of the $N$ nuclear spins within each dot have dephased through inhomogeneous broadening. This dephasing results in an effective new, random transverse polarization associated with each dot of order $\sqrt{N}$.
@@
To model these data, we consider the evolution of $\Delta_{ST}$ under the influence of nuclear spin dephasing. Assuming that at $t_{n}=0$, $\Delta_{ST}=0$, we analytically find that $\langle\Delta_{ST}^{2}\rangle\propto 1-\exp[-(t_{n}/T_{c})^{\gamma}]$. The characteristic time is closely related to the nuclear spin dephasing time, given by $T_{c}=T^{*}_{2,n}/\alpha$, with a factor $\alpha$ difference. Assuming that each nuclear spin experiences independent, Gaussian magnetic noise with a white power spectrum, $\langle\Delta_{ST}^{2}\rangle$ follows an exponential function with $\gamma=1$ and $\alpha=2$. For quasi-static nuclear spin dephasing noise, $\langle\Delta_{ST}^{2}\rangle$ follows a Gaussian function with $\gamma=2$ and $\alpha=\sqrt{2}$. By incorporating the form of $\langle\Delta_{ST}^{2}\rangle$ into the expression for the LZ transition probability $\langle P_{T}(t)\rangle\approx\frac{(2\pi)^{2}\langle\Delta_{ST}^{2}(t)\rangle}{\beta}$, we fit the experimental data. In the case of white nuclear spin dephasing noise [Fig. 3(b)], the best-fit value of $T^{*}_{2,n}$ is 4.1 ms, while for quasistatic noise [Fig. 3(c)], the best-fit value is 2.4 ms. In the figures, we also show numerical simulations with the fitted $T^{*}_{2,n}$ values. White dephasing noise, corresponding to exponential nuclear spin dephasing, appears to better describe the experimental data [Fig. 3(b)].
@@
Figure 4: Decreasing singlet-triplet relaxation. (a) Pulse sequence for measurement with and without pumping. In the case without pumping, the LZ sweep in the pump cycles is replaced by a wait pulse at $\epsilon=0$. For the probe cycles, we initialize the singlet state and pulse to different values of $\epsilon$ for a variable evolution time $t_{e}$. (b) Measured $\langle P_{T}\rangle$ for $\epsilon$ values around $\epsilon_{ST}$ in the absence of pumping. (c) Measured $\langle P_{T}\rangle$ for $\epsilon$ values around $\epsilon_{ST}$ with pumping. (d) $\langle P_{T}\rangle$ at $\epsilon_{ST}$ as a function of $t_{e}$. The experimental data (filled triangles and circles) are extracted from (b) and (c), respectively, corresponding to the $\langle P_{T}\rangle$ maximum for each $t_{e}$ value. Numerical simulations for both cases are presented.
One major potential benefit of the nuclear-spin dark state is the possibility of reduced relaxation between the singlet and triplet electronic states, due to the reduced hyperfine coupling. We discuss the possibility of reduced electron-spin dephasing in the Discussion section. In the singlet-triplet basis we use in this work, a relaxation reduction is expected because the nuclear spin operators couple to the electron raising and lowering operators, as opposed to the $\sigma^{z}$ operator (see Methods for more details on the full Hamiltonian).
@@
To test this idea, we measure the relaxation of the singlet state near $\epsilon_{ST}$ with and without DNP. In the former case, the pulse sequence starts with 64 pump cycles to form the dark state, followed by a “probe” cycle, as shown in Fig. 4(a). In the case without pumping, we use an identical pulse sequence but replace each LZ sweep in the pump cycle by a wait at $\epsilon=0$ mV, away from $\epsilon_{ST}$. In each probe cycle, with the electronic state initialized as a singlet, we pulse $\epsilon$ abruptly to a value near $\epsilon_{ST}$, wait at that $\epsilon$ for a variable time $t_{e}$, allowing the state to relax, and then abruptly pulse $\epsilon$ back to 0 mV for the measurement. Again, each probability value is obtained from over $10^{3}$ repeated measurements.
@@
Figures 4(b) and (c) show the measured $\langle P_{T}\rangle$ after the wait near $\epsilon_{ST}$ for variable time $t_{e}$ without and with pumping, respectively. Without pumping, $\langle P_{T}\rangle$ at $\epsilon_{ST}$ increases over a time-scale of about a microsecond and saturates between 0.1 and 0.2, as indicated by the yellow region in Fig. 4(b) and filled triangles in Fig. 4(d). The relatively low saturation value reflects both the relatively large voltage resolution used in this measurement, as well as the presence of detuning noise in our system. Both effects can result in a situation where the experimental detuning is not exactly $\epsilon_{ST}$, leading to a reduced transition probability.
@@
With pumping, however, $\langle P_{T}\rangle$ remains nearly zero up to at least 10 $\mu$s [Fig. 4(d)], indicating strongly suppressed singlet-triplet relaxation. In principle, $\langle P_{T}\rangle$ should remain small for up to the lifetime of the dark state, and future work will focus on confirming this hypothesis. As above, our numerical simulations agree well with our data [Fig. 4(d)]. We also note that even in the dark state, Fig. 4(c) shows that $\epsilon_{ST}$ still falls within our LZ sweep window, as discussed above.